A sweater was originally priced at $32 is on sale for $28. What is the percent of decrease rounded to the nearest whole number?
Answer by Maths68(1474)

Original price = $32
Sale price = $28
Decrease = $4
Decrease%=(Original price - Sale price)*100/(original price)
Decrease%=(32-28)*100/32
Decrease%=4*100/32
Decrease%=400/32
Decrease%=12.5%
Check
=====
12.5% of 32 = 32*(12.5/100)=400/100=4
Sale price = original price - discount
Sale price = 32-4 = 28